Shift summary, Eve 22 Dec 99

Elliott Cheu (elliott@kaon1.physics.arizona.edu)
Wed, 22 Dec 1999 23:02:56 -0700

Shift Summary for eve shift, 22 Dec 99

On Shift: E. Cheu, Wah, Sami

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.5
Hours of E799 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...):

Accounting of lost time: Starting/stopping runs, crate 8

========================================================================
Runs ended this shift:
Run 15268 - 17.4M events - BEAM_799

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: E799 datataking

CsI gains had changed due to blockhouse heater being off (since
power outage).

Crate 8 removed from readout after a flurry of errors.

Otherwise, smooth running.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15268 16:45 end run

15269 17:15 start run
22 Blockhouse heater turned on.
18:00 Request reduced to 9.5e12 to reduce 90 degree
monitor rate.
50 18:20 ktevana errors.
18:40 crate 8 removed from readout.
105 19:30 pipeline crate 6 full. Problems rebooting
crate 7. Temperature interlock had tripped off.
Reset errors -> reboot crate 7.
140 20:20 Bananas and HCC busy. Reload and kfrend.
21:30 TeV quench. Beam off for 0.5 hours.
22:00 ktevana errors.

========================================================================
Follow up Notes on Previous Shift

- updated ADGN constants.

========================================================================
Items to Follow up on Next Shift

- Be sure to install new T0's from Ashkan.
- Watch channel 968. It has a 4 count ped shift.

*** END of REPORT ***